OpenStudy (kkutie7):

Why isn't \[y(x)=e^{9x}\] the general solution of \[y'=9y\]?

jimthompson5910 (jim_thompson5910):

\[\Large y = e^{9x}\] \[\Large y \ ' = 9e^{9x}\] -------------------------------- \[\Large y \ ' = 9y\] \[\Large 9e^{9x} = 9(e^{9x})\] \[\Large 9e^{9x} = 9e^{9x} \ \ {\LARGE \color{green}{\checkmark}}\] It's definitely a solution, but is it the only solution? Are there others?

OpenStudy (wmj259):

@jim_thompson5910, nicely done!!

jimthompson5910 (jim_thompson5910):

thanks @wmj259

jimthompson5910 (jim_thompson5910):

the way to find the general solution would be to do this \[\Large y \ ' = 9y\] \[\Large \frac{dy}{dx} = 9y\] \[\Large \frac{dy}{y} = 9dx\] \[\Large \int\frac{dy}{y} = \int9dx\] \[\Large \ln(|y|) = 9x+C\] \[\Large |y| = e^{9x+C}\] \[\Large |y| = e^{9x}*e^C\] \[\Large |y| = e^C*e^{9x}\] \[\Large y = \pm e^C*e^{9x}\] \[\Large y = A*e^{9x}\] where \(\Large A = \pm e^C\). So that is the general solution which encompasses every solution

OpenStudy (kkutie7):

I got it =) once I saw your first comment I remembered there is an unknown C in 'general solutions. thanks.

jimthompson5910 (jim_thompson5910):

you're welcome